### v3

# USED: Universal Speaker Extraction and Diarization

Abstract
Speaker extraction and diarization are two enabling techniques for real-world speech applications. Speaker extraction aims to extract a target speaker’s voice from a speech mixture, while speaker diarization demarcates speech segments by speaker, annotating ‘who spoke when’. Previous studies have typically treated the two tasks independently. In practical applications, it is more meaningful to have knowledge about ‘who spoke what and when’, which is captured by the two tasks. The two tasks share a similar objective of disentangling speakers. Speaker extraction operates in the frequency domain, whereas diarization is in the temporal domain. It is logical to believe that speaker activities obtained from speaker diarization can benefit speaker extraction, while the extracted speech offers more accurate speaker activity detection than the speech mixture. In this paper, we propose a unified model called Universal Speaker Extraction and Diarization (USED) to address output inconsistency and scenario mismatch issues. It is designed to manage speech mixtures with varying overlap ratios and variable number of speakers. We show that the USED model significantly outperforms the competitive baselines for speaker extraction and diarization tasks on LibriMix and SparseLibriMix datasets. We further validate the diarization performance on CALLHOME, a dataset based on real recordings, and experimental results indicate that our model surpasses recently proposed approaches.
Index Terms:
speaker extraction, speaker diarization, multi-talker scenario, LibriMix, CALLHOME
I Introduction
Each speaker has unique voice characteristics. Speaker extraction seeks to extract speech from the target speaker with reference to his/her voiceprint [1], while speaker diarization analyzes the audio signal to determine ‘who spoke when’ [2]. As important speech processing front-ends, they are widely used in various real-world speech applications, such as speaker verification [3, 4] and speech recognition [5, 6, 7, 8].
In a multi-talker scenario, the human brain processes incoming speech signals by performing speaker extraction and diarization simultaneously. Prior studies typically treated speaker extraction [9, 10, 11] and diarization [12, 13, 14, 15, 16] as two independent tasks. However, in practical applications, such as meeting analysis and mobile recording applications, solely relying on the results from one task is insufficient since it can only provide speaking boundaries (timestamp information) or clean speech (content information). Associating extracted speech and speaker labels offers more meaningful and interpretable results as it provides a comprehensive view of ‘who spoke what and when’ [17]. Moreover, this association provides clean speech signals so that it can further facilitate the speaker-attributed automatic speech recognition task, which aims to answer ‘who spoke what’ [18, 19, 20].
Given their shared objective of disentangling speakers from a speech mixture, we believe that speaker extraction and speaker diarization complement one another. On the one hand, speaker extraction leads to a clean speech signal, therefore enhancing the effectiveness of speaker activity detection in speaker diarization. On the other hand, the information regarding the presence or absence of the target speaker resulting from speaker diarization is helpful for speaker extraction.
However, integrating speaker extraction and speaker diarization is challenging because of several critical differences between these tasks. First, their outputs are presented in very different ways. A speaker extraction system [9, 10, 11] typically generates a single speaker’s voice. In contrast, a speaker diarization system [12, 13, 14] seeks to annotate the speech activities for all speakers. Second, their application scenarios are significantly different. A speaker extraction system is usually optimized for highly overlapped speech, e.g. with a speaker overlapping ratio of near 100% [21]. Conversely, speaker diarization studies mainly handle situations with sparsely overlapped speech, as the speaker overlap ratio is around 20% in daily conversations and meetings [22, 23]. We define these two differences as the ‘output inconsistency’ and the ‘scenario mismatch’ issues, which need to be addressed when integrating them.
Most of the previous studies dealt with speaker extraction and speaker diarization either independently [13, 14, 9, 11] or jointly under some controlled conditions [24, 25], such as with a fixed number of speakers or a pre-defined overlap ratio. To date, no single-model approach has been proposed to perform speaker extraction and speaker diarization for speech mixtures with varying overlap ratios and variable number of speakers. This prompts us to look into effectively incorporating speaker extraction and speaker diarization to benefit from their interaction.
With this motivation, we propose the Universal Speaker Extraction and Diarization model (USED) to address output inconsistency and scenario mismatch. The term ‘universal’ is defined by two aspects: (a) the ability to perform for a variable number of speakers, and (b) the ability to process speech mixtures with an arbitrary overlap ratio.
To this end, we design the model by considering two aspects: data ingress and data egress. From a data ingress perspective, we design a novel embedding assignment module driven by speech references to mitigate output inconsistency naturally. This module yields two advantages. Firstly, it enables the generation of outputs for a variable number of speakers according to the speech references, thereby ensuring consistency across tasks. Secondly, it helps the model align speaker extraction and speaker diarization outputs to avoid the output permutation problem [26]. From a data egress perspective, we extend the scope of overlap ratio by a multi-task interaction module to alleviate the scenario mismatch between speaker extraction and speaker diarization. Specifically, the multi-task interaction module leverages speaker diarization results to mitigate background noise of the extracted speech and enhance the optimization of scenario-aware differentiated (SAD) loss [27] within our USED. Simultaneously, speaker extraction produces a cleaner speech signal to assist in the prediction of speaker diarization, particularly for the overlap parts.
The contributions of this paper can be summarized as follows:
  • We propose a universal solution, i.e. USED, that performs the speaker diarization and extraction jointly. The solution deals with speech mixtures with an arbitrary number of speakers and diverse overlap ratios.
  • We design an embedding assignment module to ensure consistency in the number and the order of model outputs. This aids in supporting the generation of results for a variable number of speakers and enhances the model’s robustness.
  • We design a novel multi-task interaction module with a scenario-aware differentiated loss that allows the diarization output to control whether the target speech is silenced, thereby ensuring temporal overlap consistency between the diarization and extraction outputs.
  • Our USED model outperforms competitive baseline systems of speaker extraction and speaker diarization on both the LibriMix and SparseLibriMix datasets. Additionally, we assess the diarization performance on the CALLHOME dataset, which consists of real recordings. The experimental results show that our model performs better with significantly less pre-training data.
The rest of the paper is organized as follows. Section II discusses the related work. Section III describes the proposed method. We introduce the experimental setup in Section IV. Section V presents the experimental results and analysis. Section VI concludes the study.
II Related Work
II-A Speech Separation and Speaker Extraction
Recent studies on blind speech separation have made significant progress, with the development of advanced models such as Conv-TasNet [28], Dual-Path Recurrent Neural Network (DPRNN) [29], Band-Split Recurrent Neural Network (BSRNN) [30] and efficient training algorithms such as Permutation Invariant Training (PIT) [26]. However, speech separation methods suffer from the global permutation ambiguity problem [10], which makes it hard to process long utterances. Moreover, it requires prior knowledge or estimation of the number of speakers in the speech mixture in advance, which is hard to get in real-world applications. Unlike blind speech separation that seeks to separate all speakers, speaker extraction only extracts speech for one target speaker at a time. In general, speaker extraction methods are classified into frequency-domain methods, such as SpeakerBeam [31, 32, 33] and VoiceFilter [34], and time-domain methods, e.g. SpEx+ [9] and X-SepFormer [11]. Time-domain approaches can naturally avoid the phase estimation problem, which exists in frequency-domain approaches [9]. Therefore, we develop the USED model based on the time-domain approach, SpEx+.
It is common that speech separation and speaker extraction systems in the literature are optimized for highly overlapped speech mixtures [28, 9, 11]. However, the overlap ratio in daily conversations and meetings is typically around 20% [22, 23], leading to a scenario mismatch between training on highly overlapped speech and evaluating on sparsely overlapped speech. To address such mismatch, speaker extraction for sparsely overlapped speech was recently studied [35, 36, 27]. It was suggested to minimize speech power when the target speaker is absent, while maximizing the signal-to-distortion ratio (SDR) or scale-invariant signal-to-distortion ratio (SI-SDR) [37] when the target speaker is present. This introduces another challenge since it is hard to balance the two scenario-dependent objectives, i.e., SI-SDR and power, during training. We will study a multi-task interaction module in the USED model to overcome this challenge.
II-B Speaker Diarization
Traditional clustering-based diarization methods [38, 39, 40, 41, 42] typically comprise multiple components that are trained separately. These methods implicitly assume that each speech segment is only from one speaker, which fails when multiple people speak at the same time. Two mainstream studies are attempting to solve this problem, i.e. end-to-end neural diarization (EEND) [12, 13], and target speaker voice activity detection (TS-VAD) [14]. EEND seeks to minimize the diarization error directly with permutation-free objectives [12], while TS-VAD uses the speaker embedding from the clustering results as the reference to detect the speaking status of each person. These methods have obtained good performance and set the stage for our study.
II-C Interaction Between Speech Separation and Diarization
There were studies on the interaction between speech separation and diarization [12, 43] by training speech processing models with multiple tasks. EEND-SS [25] is a joint end-to-end framework for speaker diarization and separation. It utilizes the bottleneck feature of separation as one of the inputs for diarization and is optimized using multi-task learning. This model faces challenges related to permutation and processing very long sentences. In addition, the fusion technique of EEND-SS is only applied during inference by using diarization probability directly. In TS-SEP [24], it was proposed to pre-train a model with TS-VAD objective, and subsequently finetune the model to predict the mask for separation. In other words, TS-SEP is a two-stage framework that focuses on the speech separation task. Unlike EEND-SS and TS-SEP, our USED is a universal model that can concurrently handle both tasks and address their output inconsistency and scenario mismatch issues.
Two recent works [44, 45] have explored the joint training of speech separation and speaker diarization based on unsupervised speech separation approaches. PixIT [44] employs PIT loss to integrate the speaker diarization task with MixIT [46] and draws on the best-of-both-worlds framework [47, 48] to stitch different segments based on the speaker diarization results. On the other hand, Neural FCASA [45] combines speaker diarization with neural FCA [49, 50], eliminating the need for a separate speaker diarization step to mask source activities. Both approaches jointly address the two tasks under the premise of unsupervised speech separation, focusing on long-form audio and optimizing the final separation results in terms of diarization error rate (DER) and word error rate (WER). In comparison, our USED model emphasizes the performance of supervised speech separation and speaker diarization across different overlap ratios and scenarios.
III Universal Speaker Extraction and Diarization
III-A Task Definition
Let $x$ be a multi-talker speech mixture with varying overlap ratio from 0% to 100 %, which consists of the speech signals from $l$ speakers and a noise signal $n$. The speech mixture can be represented as,
$x=\sum^{l}_{i=1}c_{i}+n$ (1)
where $l>1$, and $c_{i}$ denotes the speech signal of speaker $i$. Each speaker $i$ has a corresponding speech reference, denoted as $z_{i}$, which is estimated from the speech mixture or provided as pre-enrolled information. Note that no normalization or noise reduction techniques are applied to the raw speech signals.
Given a speech mixture and speech references, the goal of the USED model is to generate separated speech signals and speaker activity information concurrently for a variable number of speakers, ranging from $1$ to $l$. The number of speakers corresponds to the available speech references. This is motivated by scenarios where obtaining speech references for all speakers is challenging, or the target speakers are only a subset of all speakers. During inference, if there is only one target speaker, the USED model outputs a speech signal and diarization result for that speaker in a similar way to a speaker extraction model [9]. However, if there are $l$ target speakers, the model would output multiple speech signals and a diarization results concurrently, one for each speaker, just like a speech separation model does [28].
III-B System Overview
As illustrated in Fig. 1, the USED model comprises six components: a speech encoder, a speaker encoder, an embedding processing module, a separator, an extraction decoder, and a diarization decoder.
The speaker encoder converts the available speech references into speaker embeddings, that are the feature representations of the target speakers. Meanwhile, the speech encoder generates spectrum or spectrum-like feature representations from the speech mixture $x$.
Unlike TS-VAD [14], which generates diarization results for all speakers according to speaker embeddings and the speech mixture, we propose an embedding assignment module to enable custom control over the number of model’s outputs. The embedding assignment module prepares the input embeddings with three different states, active, blank, and residual, based on the speaker embeddings, which are responsible for generating results accordingly. The active state is used as a position marker for the expected speakers. The blank state serves as a position marker to guide the network in outputting a silent segment. The residual state marks the residual unexpected speakers. The blank state is motivated by Connectionist Temporal Classification (CTC) [51], which integrates blank characters to manage pauses in speech recognition and gaps between characters in Optical Character Recognition (OCR). The embedding with blank state is designed to clearly distinguish the outputs from the expected and unexpected speakers. The reason for this design is that the output from an unexpected speaker is usually a silent segment. With the help of the embedding assignment module, the USED model naturally avoids the output inconsistency stemming from the integration of speaker extraction and diarization.
The separator then aims to separate speakers at the representation level from a speech mixture and embeddings from the embedding assignment module. Finally, diarization and extraction decoders predict the corresponding results based on the output of the separator.
The USED model differs from typical speaker extraction and speaker diarization systems [12, 9] in two key aspects. Firstly, it outputs speech signals for an arbitrary number of speakers, ranging from $1$ to $l$, with the help of the embedding assignment module. Secondly, it seeks to effectively deal with speech mixtures with a wide range of overlapping ratios.
Figure 1: The overall training and inference framework of our proposed USED model. It comprises a speech encoder, a speaker encoder, an embedding assignment module, a separator, an extraction decoder and a diarization decoder. The USED model generates both the extracted speech and speaker diarization results by leveraging the speech mixture and the speech references as input. The symbols $\otimes$ and $\ominus$ refer to element-wise multiplication and frame-wise concatenation, respectively. The TCN block and mask module are denoted by rounded rectangles, which are described in detail in Fig. 2 and introduced later. Different colours are used to distinguish network layers from different components. The active, blank and residual states are assigned by Algorithm 1.
III-C Speech Encoder
Analogous to a frequency analyzer, the speech encoder aims to generate spectrum-like frame-based embedding sequences from the speech mixture and speech references. Inspired by SpEx+[9], our approach leverages a multi-scale speech encoder.
Our multi-scale speech encoder is structured as a combination of three one-dimensional Convolutional Neural Networks (Conv1D), each utilizing distinct kernel sizes to represent different scales. Subsequent to the convolutional layers, the activation function applied is the Rectified Linear Unit (ReLU).
The speech encoder produces spectrum-like frame-based embedding sequences $X$ and $Z_{i}$ for 3 scales as defined in [9] from the speech mixture $x$ and the $i$-th speech reference $z_{i}$ as
$X=[X_{1},X_{2},X_{3}]=e(x)\in\mathbb{R}^{3\times T\times C}$ (2)
$Z_{i}=[Z_{i1},Z_{i2},Z_{i3}]=e(z_{i})\in\mathbb{R}^{3\times T\times C}$ (3)
where the function $e(\cdot)$ denotes the operation performed by the multi-scale speech encoder. The set $Z=\{Z_{1},...,Z_{l}\}$ represents the embedding sequences for $l$ speakers. The input and output channel sizes of the three Conv1Ds are one and $C$, respectively. The kernel sizes of the three Conv1Ds are denoted as $L^{e}_{1}$, $L^{e}_{2}$ and $L^{e}_{3}$, while they have a common stride of $\frac{L^{e}_{1}}{2}$. The outputs $X$ and $Z_{i}$ have a shape of $3\times T\times C$, where $T$ is the number of frames. For simplicity, Fig. 1 does not explicitly depict the multi-scale structure.
(a) ResNet Block
(b) TCN Layer
Figure 2: Model structure for a ResNet block and a TCN layer. BN, PReLU, gLN and D-Conv are batch normalization, parametric ReLU, global layer normalization and dilated depth-wise separable convolution. The symbol $\oplus$ refers to element-wise addition.
Data:
$k$: Maximum number of speakers the model can process
$S_{\text{mix\_spk}}$: Set of speaker IDs for the speech mixture
$S_{\text{all\_spk}}$: Set of speaker IDs for the whole dataset
$V_{\text{em}}$: Empty embedding
$V_{\text{res}}$: Residual embedding
$p_{\text{sel}}$: Probability for selecting present speaker
$p_{\text{em}}$: Probability to use empty embedding as input
$p_{\text{res}}$: Probability to use residual embedding as input
// List of embeddings with active and blank state
1 $I_{active},I_{blank}\leftarrow[$ ],[],[] , [ ]]]];
// Number of embeddings
2 $count\leftarrow 0$;
/* The First Step for Active State */
3 for $s_{\text{mix}}\in S_{\text{mix\_spk}}$ do
4       if $s_{\text{mix}}$ is a present speaker then
5             $p\leftarrow$ random_uniform$(0,1)$;
             // If true, then marking embedding as active state
6             if $p>p_{\text{sel}}$ then
7                   $E_{count}\leftarrow V_{s_{\text{mix}}}$;
8                   $I_{active}$.add($E_{count}$);
9                   $count\leftarrow count+1$;
10
11             end if
12
13       end if
14
15 end for
16
/* The Second Step for Blank State */
17 while $count\leq k$ do
18       $p\leftarrow$ random_uniform$(0,1)$;
19      if $p>p_{\text{em}}$ then
             // Randomly select one speaker from the set of speakers that includes $S_{\text{all\_spk}}$ but excluding $S_{\text{mix\_spk}}$
20             $s_{\text{ot}}\leftarrow$random_sample$(S_{\text{all\_spk}}\setminus S_{\text{mix\_spk}})$;
21             $E_{count}\leftarrow V_{s_{\text{ot}}}$;
22
23      else
24             $E_{count}\leftarrow V_{\text{em}}$;
25
26       end if
27      $I_{blank}$.add($E_{count}$);
28       $count\leftarrow count+1$;
29
30 end while
/* The Third Step for Residual State */
31 $p\leftarrow$ random_uniform$(0,1)$;
32 if $p>p_{\text{res}}$ or $I_{active}.\text{length}\neq$ number of present speakers then
33       $E_{k+1}\leftarrow V_{\text{res}}$;
34
35else
36       $E_{k+1}\leftarrow V_{\text{em}}$;
37
38 end if
Output: $I_{active}$, $I_{blank}$, $E_{k+1}$
Algorithm 1 Assigning States for Embeddings
III-D Speaker Encoder
The speaker encoder is designed to extract speaker embeddings from the corresponding enrolled speech. Specifically, we employ a speaker encoder to extract speaker embeddings, denoted as $V_{1},...,V_{l}$, from the embedding sequences $Z$. These speaker embeddings capture the distinctive characteristics of each speaker. After layer normalization (LN), we leverage a CNN layer with a kernel size of $1\times 1$ ($1\times 1$ CNN) on the embedding sequence $Z_{i}$ for speaker $i$. This is subsequently followed by $N^{r}$ residual network (ResNet) blocks [52]. The input and output channel sizes of this $1\times 1$ CNN are $3C$ and $C$, respectively. Finally, an additional $1\times 1$ CNN, combined with a mean pooling operation, generates a speaker embedding $V_{i}$ for speaker $i$ with an embedding dimension $D_{\text{spk}}$. After that, the speaker encoder predicts speaker identities, denoted as $\hat{y}^{\text{spk}}_{i}$, through a linear layer followed by a softmax (SM) activation function for speaker $i$.
The detailed architecture of the ResNet we used is shown in Fig. 2a. A ResNet block comprises two $1\times 1$ CNNs and a max-pooling layer. Each $1\times 1$ CNN within the ResNet block is accompanied by batch normalization (BN) and parametric ReLU (PReLU) for normalization and non-linear transformation. A skip connection adds the input to the output obtained after the second BN layer. Lastly, the max-pooling layer is responsible for adjusting the length of the output embedding.
We choose ResNet [52] because it has become a prominent architecture in speaker verification systems [53, 54, 55, 56] and is widely adopted in speaker extraction models [9, 57]. Its key advantage lies in the residual connections, which link frame-level layers and mitigate the vanishing gradient problem, facilitating faster convergence during backpropagation [52]. These residual connections also enable deeper neural network construction, allowing ResNet-based models to outperform regular CNNs. As a result, using ResNet as the backbone for speaker encoders leads to more robust and higher-quality speaker embeddings, making it a natural choice for this task.
III-E Embedding Assignment Module
The embedding assignment module aims to prepare embeddings with different states for the separator. We set embeddings with different states: active, blank, and residual. The details of the embedding assignment module are shown in Algorithm 1.
III-E1 Active State
The embedding with active state is designed to guide the network to generate outputs for the expected speakers. Based on the embedding with active state, the whole speaker extraction and diarization process is driven by the speaker encoder that takes the speech references of some present speakers as input. Specifically, to enable our model to achieve the correct output for an arbitrary number of speakers, i.e., from 1 to $l$, we randomly assign the active state to the embeddings of present speakers by a hyperparameter $p_{\text{sel}}$ representing probability, as shown in the first step of Algorithm 1.
III-E2 Blank State
The embeddings with blank state are primarily responsible for driving the network to produce silent segments for the unexpected speakers, which is critical in ensuring our network’s robustness. Unlike active state, which requires a unique embedding from the speaker encoder, the embeddings with blank state alternate between two embeddings via a threshold $p_{\text{em}}$ to guide the network: an inactive speaker embedding and a learnable empty embedding. The inactive speaker embedding is derived from a speaker who is not present in the speech mixture. This design ensures that our USED model can customize the number of outputs by inputting the learnable empty embedding during the inference stage.
III-E3 Residual State
As an additional state, the targets corresponding to the embedding with residual state are the cumulative outputs for speakers that are present but are not selected at the first step. This design allows our USED model to explore the interaction among all speakers in the speech mixture. Specifically, we also employ a learnable embedding $V_{\text{res}}$ to teach the network to learn the corresponding cumulative outputs for the residuals. When speaker embeddings of all speakers have been considered as the active state, the target corresponding to residual state should be left without any speakers. In such cases, the choice between $V_{\text{res}}$ and $V_{\text{em}}$ is determined through a specified threshold $p_{\text{res}}$. Our preliminary experiments demonstrate that this strategy enhances robustness during the inference phase.
Overall, our USED model has $(k+1)$ input embeddings, where the first $k$ embeddings with active and blank states support the predictions of at most $k$ speakers and the $(k+1)$-th embedding $E_{k+1}$ is with the residual state. Finally, we shuffle all the embeddings except $E_{k+1}$ to guarantee the model’s insensitivity to the order of embeddings. Through the designed embedding processing mechanism, our USED model can predict the outputs from a custom number of speakers. Additionally, the model’s robustness is enhanced by employing a random selection strategy for speakers when assigning blank state.
III-F Separator
The separator is designed to separate speakers based on the embeddings from the embedding assignment module. As shown in Fig. 1, the embedding sequence $X$ firstly undergoes a series of operations, including a layer normalization followed by $1\times 1$ CNN layer, wherein the input channel size is set to $3C$ and the output channel size is $C$. Then, we introduce $N^{t}$ Temporal Convolutional Network (TCN) blocks [28] to generate output representations based on the embeddings and speech mixture.
We use TCN as the backbone of our model because they have been widely adopted in speech separation and speaker extraction tasks[28, 9]. The USED model is built on the SpEx+ model, which also utilizes TCN. This choice ensures consistency and enhances performance. In addition, TCN is commonly used for time-domain approaches, avoiding the phase estimation issues found in time-frequency models. Finally, compared to methods like DPRNN [29] and BSRNN [30], Conv-TasNet [28] offers lower computational complexity, reduced memory requirements, and faster training and inference.
Representations are then concatenated and down-sampled via a Conv1D layer with input channel size $(k+1)\times C$ and output channel size $C$. Finally, the processed representation is forwarded to another $N^{t}$ TCN blocks to generate representations $\{R_{1},...,R_{{}_{N^{t}}}\}$.
Each TCN block comprises a set of $N^{b}$ TCN layers. The dilated depth-wise separable convolution (D-Conv) shown in Fig. 2b has an exponential growth dilation factor $2^{b}$, where $b\in\{0,...,N^{b}-1\}$. In the initial TCN layer within each block, the input channel size for the $1\times 1$ CNN is $D_{\text{spk}}+C$.
III-G Diarization Decoder
The diarization decoder aims to predict the frame-level activity probabilities for each speaker from the $N^{t}$ TCN blocks’ outputs $\{R_{1},...,R_{{}_{N^{t}}}\}$ in the separator. Because speaker extraction and diarization require different time resolutions, we employ a Conv1D layer to down-sample the outputs from the TCN block, with input and output channel sizes of $C$, a kernel size of $L^{\text{diar}}$, and a stride of $L^{\text{diar}}/2$. Subsequently, we use a linear layer followed by a softmax operation to compute the probabilities, denoted as $D^{j}=\{D^{j}_{1},...,D^{j}_{k+1}\}$, which represent the speech activities of the corresponding embeddings. Finally, we get a set of diarization results from each TCN block, denoted as $\{D^{1},...,D^{N^{t}}\}$. During inference, we only use the diarization result $D^{N^{t}}$ from the final TCN block as the results to calculate the evaluation metric.
III-H Extraction Decoder
The extraction decoder is designed to estimate masks for each speaker and reconstruct the corresponding signals. The mask module of the extraction decoder comprises a Conv1D layer and ReLU activation to generate masks $\{M_{1},...,M_{k+1}\}$, where $M_{i}\in\mathbb{R}^{T\times C}$. We then obtain the modulated responses $i$ by element-wise multiplication of the mask $M_{i}$ and the representations $X$ from the speech mixture.
For signal reconstruction, we employ a multi-scale decoder like speech encoder, which consists of three transposed convolutional networks (ConvTran1D), each with a kernel size of $L^{\text{e}}_{1}$, $L^{\text{e}}_{2}$ and $L^{\text{e}}_{3}$, and a common stride, $L^{\text{e}}_{1}/2$. It’s used to reconstruct the time-domain signals, $\tilde{s}=\{\tilde{s}_{1},...,\tilde{s}_{k+1}\}$, where $\tilde{s}_{i}=\{\tilde{s}^{1}_{i},\tilde{s}^{2}_{i},\tilde{s}^{3}_{i}\}$ is obtained from three ConvTran1D layers.
As the USED model generates both the extracted speech and speaker diarization results, we propose an interaction module (IM) to use speaker diarization results to refine the generated waveforms. To avoid the gradient from the extraction task having an effect on the diarization task, we first create a clone of $D^{N^{t}}=\{D^{N^{t}}_{1},...,D^{N^{t}}_{(k+1)}\}$ without gradient. Subsequently, we employ interpolation techniques on the diarization output probabilities to ensure length-based uniformity between the two tasks. Following this alignment, a layer comprising a Conv1D layer and ReLU activation is employed to further process the output probabilities and obtain outputs $r_{i}$, ensuring that the values at each step are not constrained to be less than 1. The Conv1D layer possesses an input channel size of 1, an output channel size of 1, and a kernel size of $L^{\text{im}}$.
Finally, the model generates the extracted speeches, $\{\hat{s}_{1},...,\hat{s}_{k+1}\}$, through element-wise multiplication of $\{r_{1},...,r_{k+1}\}$ and the output $\tilde{s}$ from the extraction module separately as
$\hat{s}_{i}=\{\hat{s}^{1}_{i},\hat{s}^{2}_{i},\hat{s}^{3}_{i}\}$ (4)
where
$\hat{s}^{j}_{i}=\tilde{s}^{j}_{i}\otimes r_{i}$ (5)
It should be noted that, during inference, only $\tilde{s}^{1}_{1},...,\tilde{s}^{1}_{k}$ originating from the first ConvTran1D are utilized for the evaluation of metrics.
III-I Loss Function
III-I1 Overall Loss Function
The loss of the USED model can be formulated as follows:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{ext}}+\lambda_{2}\mathcal{L}_{\text{diar}}+\lambda_{3}\mathcal{L}_{\text{spk}}$ (6)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the weights for speaker extraction loss, speaker diarization loss, and speaker classification loss, respectively.
Figure 3: Illustration of scenario-aware differentiated loss. The speech mixture is segmented according to the four scenarios (QQ, QS, SS and SQ).
III-I2 Speaker Extraction
We adopt the SAD loss proposed in [27] to address sparsely overlapped speech as shown in Fig. 3. Scenarios are classified into four distinct classes: QQ, QS, SS, and SQ. In the QQ scenario, both the target speaker and interference speakers are in a quiet state. The QS scenario represents a situation where the target speaker is quiet while the interference speakers are speaking. Conversely, in the SQ scenario, the target speaker is speaking while the interference speakers are quiet. Lastly, the SS scenario corresponds to both the target speaker and interference speakers being in a speaking state. For QS and QQ, we use power as the objective for extracted speech $\hat{s}$,
$\mathcal{L}_{\text{E}}=\sum^{3}_{j=1}10\mu_{j}\log_{10}(\frac{\|\hat{s}^{j}\|^{2}}{T_{\text{se}}}+\epsilon)$ (7)
where $T_{\text{se}}$ is the duration of $\hat{s}^{j}$ in seconds, $[\mu_{1},\mu_{2},\mu_{3}]$ are the weights for different ConvTran1D layers and $\epsilon$ is set to $10^{-6}$. For SS and SQ, we compute the SI-SDR loss between the target speech $s$ and extracted speech $\hat{s}$ as $\mathcal{L}_{\text{S}}$,
$\mathcal{L}_{\text{S}}=\sum^{3}_{j=1}-10\mu_{j}\log_{10}\frac{\|\frac{<\hat{s}^{j},s>s}{\|s\|^{2}+\epsilon}\|^{2}}{\|\hat{s}^{j}-\frac{<\hat{s}^{j},s>s}{\|s\|^{2}+\epsilon}\|^{2}+\epsilon}+\epsilon$ (8)
SAD loss is used to calculate the loss between the extracted speeches $\{\hat{s}_{1},..,\hat{s}_{k+1}\}$ generated from the multi-task interaction module and target speeches $\{s_{1},...,s_{k+1}\}$. For clarity, we assume the first $m$ target speeches correspond to the embeddings with active state and are the same as $\{c_{1},...,c_{m}\}$ in Eq. (1). $\{s_{m+1},...,s_{k}\}$ are zero vectors which present silence as the target of embeddings with blank state. $s_{k+1}$ is the target speech for residual prediction. In practice, the order is shuffled as introduced in section III-E. Each pair of extracted speech and target speech are firstly segmented according to the four scenarios. Finally, the total speaker extraction loss, i.e. $\mathcal{L}_{\text{ext}}$, is as follows:
$\mathcal{L}_{\text{ext}}=\frac{1}{k+1}(\sum_{i=1}^{k+1}\alpha\mathcal{L}_{\text{E}_{i}}^{QQ}+\beta\mathcal{L}_{\text{E}_{i}}^{QS}+\gamma\mathcal{L}_{\text{S}_{i}}^{SS}+\delta\mathcal{L}_{\text{S}_{i}}^{SQ})$ (9)
where $k+1$ is the total number of outputs and $\alpha$, $\beta$, $\gamma$ and $\delta$ are the weights for different scenarios.
III-I3 Speaker Diarization
To optimize the model for the diarization task, we apply the binary cross-entropy loss (BCE), denoted as $\mathcal{L}_{\text{diar}}^{j}$ on the output of the $N^{t}$ TCN blocks,
$\mathcal{L}_{\text{diar}}^{j}=\frac{1}{k+1}\sum_{i=1}^{k+1}BCE(y^{\text{diar}}_{i},D_{i}^{j})$ (10)
where $y^{\text{diar}}_{i}$ is the ground-truth label for the $i$-th output. Then, the entire speaker diarization loss can be represented as:
$\mathcal{L}_{\text{diar}}=\sum_{j=1}^{N^{t}}\mathcal{L}^{j}_{\text{diar}}$ (11)
III-I4 Speaker Classification
The speaker encoder is optimized through a cross-entropy loss (CE), denoted as $\mathcal{L}_{\text{spk}}$, for speaker classification during the training process, as
$\mathcal{L}_{\text{spk}}=\frac{1}{l}\sum_{i=1}^{l}CE(y^{\text{spk}}_{i},\hat{y}^{\text{spk}}_{i})$ (12)
where $y^{\text{spk}}_{i}$ and $\hat{y}^{\text{spk}}_{i}$ are the ground truth and predicted speaker ids for speaker $i$, respectively.
IV Experimental Setup
IV-A Datasets
We conduct comprehensive evaluations of our model using diverse datasets, including LibriMix and SparseLibriMix, which are simulated, and CALLHOME, which consists of real recordings.
IV-A1 LibriMix
The first dataset, known as LibriMix [21], generates samples derived from LibriSpeech [58] train-clean-100 and test-clean datasets, as well as WHAM! [59], all with a 16kHz sampling rate. LibriMix has been widely used for speech separation [60, 61], speaker extraction [62] and speaker diarization tasks[61, 25]. We merge the Libri2Mix 100h and Libri3Mix 100h datasets, resulting in utterances that may contain either 2 or 3 speakers.
The models are evaluated on both min and max modes. The min mode consists of highly overlapped speech segments, while a larger portion of the max mode consists of non-overlapped speech segments. To ensure compatibility with the speaker extraction task, we have slightly adjusted the data split for training and validation  [^1] https://github.com/msinanyildirim/USED-splits, as previously done in related work [62]. Specifically, we follow scripts  [^2] https://github.com/gemengtju/L-SpEx/tree/main/data to prepare speech references for the Libri2Mix dataset and randomly select speech references for each speech mixture in the Libri3Mix dataset.
IV-A2 SparseLibriMix
The second dataset employed in our evaluation is the test set from SparseLibriMix [21], which utilizes WHAM! [59] to simulate a noisy acoustic environment. This dataset covers more realistic mixture scenarios, varying from an overlap ratio of 0 to 1.0. It contains 1,000 utterances which have 2 or 3 speakers.
IV-A3 CALLHOME
We use the CALLHOME dataset [63] to evaluate the diarization task with real recordings. The CALLHOME dataset is partitioned into two parts based on the Kaldi recipe  [^3] https://github.com/kaldi-asr/kaldi/tree/master/egs/callhome_diarization/v2. The first part, known as Part 1, is utilized for model adaptation, whereas the second part, referred to as Part 2, is used for evaluation. The CALLHOME dataset consists of telephone-channel recordings with 8k sample rate. Besides CALLHOME Part 1, we further add Libri2Mix 100h, Libri3Mix 100h, Libri2Mix 360h and Libri3Mix 360h, which are generated from LibriSpeech [58] train-clean-100 and train-clean-360, as the pre-training dataset. In total, our dataset encompasses approximately 452 hours of data. We then finetune the model only with CALLHOME Part 1 and evaluate on CALLHOME Part 2. During inference, we follow the evaluation pipeline of TS-VAD to evaluate our model so that it does not require pre-enrolled speech. Specifically, we extract the speech references based on the diarization result of spectral clustering and then do inference for USED models. The speaker embedding for spectral clustering is extracted from the ECAPA-TDNN model [64] trained on VoxCeleb2 [65]  [^4] https://github.com/TaoRuijie/ECAPA-TDNN.
IV-B Baseline Configuration
Two of the baseline models, SpEx+ [9], and TS-VAD [14], are implemented by ourselves. During the training phase, we segment the utterances into chunks, each of which spans 4 seconds and is shifted every 2 seconds.
For the SpEx+ baseline, we use the same training configuration as the USED model described in section IV-C. To align the model complexity with that of the USED model, we configure the SpEx+ baseline with a total of 8 TCN blocks so that it has a similar number of parameters to the USED model. It is worth noting that the SpEx+ baseline can be regarded as a specialized version of the USED model, as it has only one embedding with active state as input without including a diarization module and SAD loss. Our metric for selecting the best checkpoint for the SpEx+ model is SI-SDR.
As for the TS-VAD baseline, we follow [14] to design the model, with several notable modifications. Firstly, we substitute the bidirectional LSTM in the original TS-VAD model with four transformer layers, with sine-cosine positional encoding as input. The model architecture entails two layers dedicated to processing inputs for each speaker individually, followed by a Conv1D layer for down-sampling. Another two layers are used to combine the information from all speakers. For the transformer, we set the embedding dimension to 384 and employ four attention heads. Second, we use a pre-trained speaker encoder, ECAPA-TDNN [64], to extract speaker embeddings. This speaker encoder also takes on the role of replacing the CNN encoder to extract time-level representations from the speech mixture. Finally, we use a similar strategy as stated in section III-E, which randomly selects speaker embeddings from the whole dataset. $p_{em}$ is set to be 0.3. The model is optimized using Adam with batch size 64 for 40k steps. The speech encoder is kept fixed for the first 4,000 steps, and the learning rate follows a polynomial decay schedule with an initial value of $2\times 10^{-4}$, which has a warm-up phase for the first 10% of steps. The DER metric is our criterion for selecting the best checkpoint for the TS-VAD model.
IV-C USED Model Configuration
The maximum number of speakers $k$ is 3 for evaluation on the LibriMix and SparseLibriMix datasets and in the assessment on the CALLHOME dataset. For the multi-scale speech encoder, $C$ is 256 and the kernel sizes, $L^{e}_{1}$, $L^{e}_{2}$ and $L^{e}_{3}$, are 20, 80 and 160. The number of ResNet blocks $N^{r}$ in the speaker encoder is set to 4, and the dimension of the speaker embedding $D_{\text{spk}}$ is 256. Regarding the separator module, we set $N^{t}$ to 3 and $N^{b}$ to 8. For the diarization decoder, the CNN layer has a kernel size $L^{\text{diar}}$ of 32 and stride $L^{\text{diar}}/2$ of 16. The mask module of the extraction decoder consists of a CNN with a kernel size of 1 and stride 1, and ReLU activation. For the multi-task interaction module, the kernel size $L^{\text{im}}$ of the CNN layer is 16. The loss weights for outputs from different ConvTran1D layers, $\mu_{1}$, $\mu_{2}$ and $\mu_{3}$ are 0.8, 0.1 and 0.1 by following [9]. We set $0.001$ for $\alpha$ and $\beta$, and $1.0$ for $\gamma$ and $\delta$, which are the loss weights in Eq. (9). $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 1.0. $p_{\text{em}}$ is set to 0.3. $p_{\text{sel}}$ and $p_{\text{res}}$ are set to 0.5 and 0.9, respectively, when residual state is applied. The USED model without residual state and with $p_{sel}=0$ is called as “USED-F(ix)” since it always gets speech references of present speakers during training.
We optimize the model with Adam on 4 GPUs for 100k steps, corresponding to approximately 11 epochs. The utterances are split into chunks with a size of 4 seconds and a shift of 2 seconds. We set the maximum tokens to 260k, which results in processing around 16 seconds of audio for each batch. The learning rate is set to $31e-$ 3 for all experiments, which is warmed up for the first 10% steps and decayed polynomially for the remaining steps. The overall loss value is our primary metric for selecting the best checkpoint for the USED model.
IV-D Evaluation Metrics
We use diarization error rate (DER (%)), including speaker confusion (SC (%)), false alarm (FA (%)), and missed detection (MS (%)) to evaluate the speaker diarization performance. Collar tolerance and median filtering are set to 0 seconds and 11 frames for the LibriMix and SparseLibriMix datasets, and 0.25 seconds and 11 frames for the CALLHOME datasets. We report SI-SDR improvement (SI-SDRi (dB)), Power (dB/s), SDR improvement (SDRi (dB)), Short-Time Objective Intelligibility (STOI) [66] and Perceptual Evaluation of Speech Quality (PESQ) [67] for speaker extraction performance. In terms of complexity analysis, we employ Multiply-Accumulate Operations (MACs (G)) to assess the computational cost of the model and perform a calculation of the model’s parameter count.
TABLE I: Performance on the LibriMix dataset for max mode. The results of baseline systems are obtained by our own implementations.
Speaker Diarization
Model Overall Performance SS SQ&QS QQ
[…]
wav2vec 2.0 Base [68] 7.62 2.28 4.82 0.52 - - -
HuBERT Base [69] 7.56 2.40 4.81 0.35 - - -
TS-VAD [14] 7.28 3.61 2.78 0.89 5.71 11.63 0.08
USED-F 4.75 2.18 2.16 0.42 3.21 9.41 0.05
USED 5.20 2.68 2.08 0.43 3.56 10.36 0.04
Speaker Extraction
Model Overall Performance SS SQ QS&QQ
[…]
SpEx+ [9] 9.06 10.11 0.776 1.39 8.32 5.17 55.60
USED-F 12.70 13.22 0.844 1.46 12.00 9.17 -24.00
USED 12.22 12.69 0.836 1.43 11.63 9.30 -14.15
V Experimental Results and Analysis
V-A Results on the LibriMix Dataset
Tables I and II show the results on the LibriMix dataset with max mode and min mode, respectively. As indicated before, min mode mainly contains highly overlapped speech, while a larger proportion of max mode is non-overlapped speech compared to min mode. We provide detailed results under different scenarios for the max mode as shown in Table I. Please note that we use the average duration length that anyone is speaking in seconds as the metric of speaker diarization for the QQ scenario.
Our analysis involves comparisons with other popular systems from the literature, including TS-VAD [14], EEND-EDA [13], HuBERT Base [69] and wav2vec 2.0 Base [68] for speaker diarization, SpEx+ [9], and Conv-TasNet [28] for speaker extraction and speech separation, and EEND-SS [25] for joint optimization. For HuBERT Base and wav2vec 2.0 Base, we firstly follow the SUPERB diarization task [61] to train two downstream models for two speakers and three speakers, respectively. Subsequently, we combine the output results of these two models for evaluation.
The proposed USED model demonstrates superior or comparable performance compared to the baseline systems, as indicated by their respective evaluation metrics in both the min and max modes. For speaker diarization, our models exhibit remarkable improvements, with a relative reduction in DER of at least 13% when compared to the prior systems. For speaker extraction, our models achieve a much lower value in terms of power compared to other baseline systems under QQ and QS scenarios in the max mode. This phenomenon implies that speaker diarization and SAD loss effectively mitigate background noise during silent intervals, resulting in enhanced performance.
[…]
Speaker Diarization
Model DER $\downarrow$ MS $\downarrow$ FA $\downarrow$ SC $\downarrow$
EEND-EDA [25] 10.16 - - -
TS-VAD †[14] 5.30 2.32 2.83 0.15
EEND-SS [25] 6.27 - - -
   + LMF 6.04 - - -
USED-F 4.64 1.99 2.55 0.10
USED 4.57 1.94 2.54 0.09
Speaker Extraction / Speech Separation
Model SI-SDRi $\uparrow$ SDRi $\uparrow$ STOI $\uparrow$ PESQ $\uparrow$
Conv-TasNet [25] 7.66 8.71 0.756 -
SpEx+ †[9] 8.76 10.09 0.772 1.44
EEND-SS [25] 9.31 7.50 0.760 -
   + Fusion 9.38 7.59 0.760 -
   + LMF 8.83 9.72 0.767 -
   + LMF + Fusion 8.87 9.77 0.767 -
USED-F 12.24 12.83 0.841 1.54
USED 11.98 12.50 0.834 1.44
V-B Results on SparseLibriMix
In addition to the LibriMix dataset, we are also interested in assessing the performance of our model across a range of overlap ratios. Therefore, we conduct an evaluation on SparseLibriMix, as illustrated in Fig. 4. We select TS-VAD and SpEx+ models as the baseline systems. The model trained on the max mode of Libri2Mix and Libri3Mix is evaluated on SparseLibriMix with 2 and 3 speakers. The proposed USED models outperform each task baseline, confirming our models’ effectiveness for variable values of overlap ratio.
(a) Speaker Diarization
(b) Speaker Extraction
Figure 4: Performance on the SparseLibriMix for overlap ratio from 0% to 100%. The Left and right bar charts are the results of speaker diarization and speaker extraction tasks, respectively.
V-C Results on CALLHOME
To compare the performance of diarization on real data, we evaluate our model on the CALLHOME dataset. The results are summarized in Table III, compared with several speaker diarization methods, including two clustering-based methods, AHC clustering [13] and spectral clustering [41], PixIT [44], end-to-end speaker diarization models, EEND-EDA [13], SC-EEND [70] and AED-EEND [71], and the TS-VAD model [14]. Most of those methods use the simulation data from Swichboard-2, Switchboard Cellular, and NIST Speaker Recognition Evaluation [13], which have a variable number of speakers, ranging from one to five. The total number of hours for the pre-training data is around 15,516 hours for those methods. In contrast, we use much less simulation data, which is from LibriMix and only have 2 or 3 speakers. This results in around 452 hours of pre-training data.
Our USED model achieves around 15% relative improvement on overall DER compared to the TS-VAD baseline and performs better with much less pre-training data compared to other methods. We observe that our USED model is more robust in terms of the number of speakers. Results show that although our model only uses simulation data from LibriMix, which has a speech mixture with 2 or 3 speakers, the model still achieves better DER results when the number of speakers is larger than 3. We believe this gain is from the design of the embedding assignment module.
[…]
Model Pre-training Data Number of Speakers
2 3 4 5 6 all
AHC clustering [13] - 15.45 18.01 22.68 31.40 34.27 19.43
spectral clustering [41] - 16.05 18.77 22.05 30.46 36.85 19.83
PixIT †[44] 79h 17.02 26.03 27.75 36.38 40.77 24.41
EEND-EDA [13] 15,516h 8.50 13.24 21.46 33.16 40.29 15.29
SC-EEND [70] 15,516h 9.57 14.00 21.14 31.07 37.06 15.75
AED-EEND [71] 15,516h 6.96 12.56 18.26 34.32 44.52 14.22
TS-VAD $\ddagger$ [14] 452h 12.95 14.57 20.26 27.36 32.66 15.51
USED-F 452h 9.09 12.76 14.67 26.96 26.71 13.16
USED 452h 10.23 12.78 14.68 32.02 25.10 13.51
TABLE IV: Model performance on the LbriMix dataset for max mode when setting different values of $m$. RS stands for residual state.
$m$ Model SI-SDRi $\uparrow$ DER $\downarrow$
1 SpEx+ [9] 9.06 -
USED-F 5.71 20.30
USED 9.73 11.13
   - w/o RS 9.87 11.96
   - $p_{\text{res}}=1.0$ 10.23 10.85
2 USED-F 9.55 11.27
USED 11.02 9.42
   - w/o RS 10.96 10.98
   - $p_{\text{res}}=1.0$ 11.08 10.14
all TS-VAD [14] - 7.28
USED-F 12.70 4.75
USED 12.22 5.20
   - w/o RS 11.80 9.46
   - $p_{\text{res}}=1.0$ 11.76 10.38
V-D Investigation of USED Models for Generating Results of Different Numbers of Speakers During Inference
In this section, we compare the USED-F model and the USED model with baseline systems when generating results for different numbers of speakers during inference, as shown in Table IV. As introduced in Section III-I2, $m$ is the number of embeddings with an active state as the input of the USED model. During inference, $m$ also represents the number of speakers the USED model is required to predict. When $m$ is 1, USED models perform speaker extraction like SpEx+. Under this condition, USED models and SpEx+ have the same input data for a fair comparison. When $m$ is all, it means we provide all speech references for present speakers, and the embedding is $V_{\text{em}}$ for residual state. Otherwise, we use $V_{\text{res}}$ as the embedding for residual state.
First, USED models consistently perform much better than the SpEx+ baseline when $m$ is 1, except for the USED-F model. This indicates that even if the USED model only gets one speech reference, it can still outperform the SpEx+ model. When $m$ is 2 or all, the performance of USED in speaker extraction significantly surpasses that of SpEx+. The experimental results show that the USED model demonstrates superior efficiency and performance in speaker extraction tasks, particularly in multi-speaker scenarios. Unlike SpEx+, which requires multiple inferences for each speaker, USED performs inference just once, reducing redundant computations like those in the speech encoder. It can also simultaneously utilize the embeddings of multiple speakers, resulting in more discriminative results, whereas SpEx+ processes one speaker at a time. As the number of speakers increases, USED’s performance advantage becomes even more significant.
Second, when extracting results for all speakers, the USED model outperforms TS-VAD, as shown in the experimental results. In addition, the TS-VAD model can only extract results for all speakers simultaneously, while the USED model provides flexibility to extract results for a specific speaker or multiple speakers, which is advantageous in scenarios where obtaining embeddings for all speakers at once is not feasible. Furthermore, the USED model does not require a pre-trained speaker encoder, unlike TS-VAD, which typically relies on one.
Third, we observe that the USED model without residual state achieves improvement compared to the USED-F model on both speaker diarization and extraction tasks if $m$ equals 1 or 2. It gets a larger improvement when applying residual state with $p_{\text{res}}$ of 1.0. However, when $m$ is all, the performance of these two models degrades significantly on the speaker diarization task compared to the USED-F model with the DER increasing from 4.75% to 10.38% and 9.46%, respectively. This degradation can be avoided by setting $p_{\text{res}}$ to 0.9 during training. In this way, the USED model can achieve a competitive performance compared to that of the USED-F model if $m$ is all and still significantly outperforms baseline systems when $m$ is 1 or 2.
The performance gap between USED-F and USED models when $m=\text{all}$ is due to the fact that the input to the USED-F model consistently includes speaker embeddings for all speakers during training. As a result, USED-F is designed to predict for all speakers simultaneously, making it specifically optimized for scenarios where $m=\text{all}$. However, its performance is less robust in cases where $m\neq\text{all}$ (e.g., $m=1$ or $m=2$), leading to lower performance compared to the USED model in these situations.
TABLE V: USED-F model performance under different scenarios when using different loss weights for each scenario of SAD loss with or without Multi-Task interaction module. IM means multi-task interaction module. The results of the first line are from the standard setting.
$\alpha$ $\beta$ $\gamma$ $\delta$ IM Speaker Extraction Speaker Diarization
[…]
All SS SQ QS & QQ
0.001 0.001 1.0 1.0 12.70 12.00 9.17 -24.00 4.75
0 0 1.0 1.0 12.47 12.02 9.58 55.68 4.80
0.001 0.001 1.0 1.0 12.71 12.00 9.11 -1.93 4.65
0.01 0.01 1.0 1.0 12.50 11.65 8.68 -45.16 4.77
1.0 1.0 1.0 1.0 11.54 10.48 7.42 -79.21 5.57
TABLE VI: Ablation study of the USED model configured for a single task (Only SD or Only SE) on the LibriMix dataset for max mode.
Speaker Diarization
Model All SS SQ&QS QQ
[…]
USED-F 4.75 3.21 9.41 0.05
   - Only SD 6.43 4.72 11.76 0.04
Speaker Extraction
Model All SS SQ QS&QQ
[…]
USED-F 12.70 12.00 9.17 -24.00
   - Only SE 12.46 11.71 9.03 20.84
V-E Ablation Study: Assessing Single-Task Training for the USED Model
We then compare USED models trained on single task to investigate the impact of each task as illustrated in Table VI. Without loss of generality, we use the USED-F model for comparison. We present the results obtained by setting $\lambda_{1}$ or $\lambda_{2}$ to 0, referred to as Only SD (Speaker Diarization) and Only SE (Speaker Extraction), respectively.
The USED-F model achieves a notable reduction of DER with a relative improvement of around 32% in the SS scenario and 20% in the SQ&QS scenarios. The larger improvement in the SS scenario suggests that speaker extraction indeed plays a pivotal role in enhancing the diarization module’s ability to handle overlapping speech segments. Furthermore, SI-SDRi increases from 12.46 dB to 12.70 dB for overall performance when compared to the USED-F model with only SE, which may be mainly from the improvement under the QQ scenario where the power reduces from 20.84 dB/s to -24.00 dB/s. This indicates that the speaker diarization task helps surpass the background noise for the speaker extraction task.
V-F Investigation of the Relationship between the Multi-Task Interaction Module and the SAD Loss
TABLE VII: Model Comparison in Terms of Parameters and MACs for the TS-VAD model, the SpEx+ model and the USED model. “Params” refers to the number of model parameters, which is counted in millions (M). †The MACs of SpEx+ is only for extracting one speaker’s result.
Model Params MACs Total MACs
TS-VAD [14] 39.50 27.21 158.01
SpEx+ †[9] 16.35 43.60
USED-F 23.12 96.91 96.91
USED 23.65 119.54 119.54
In this section, we analyze the performance of the USED-F model using different weights for each scenario of SAD loss with or without the multi-task interaction module as illustrated in Table V. For the experiments without the multi-task interaction module, results demonstrate that when increasing the weights $\alpha$ and $\beta$ for QQ and QS scenarios, we can get a lower power value, but the SI-SDRi are reduced from 12.02 dB to 10.48 dB for the SS scenario and from 9.58 dB to 7.42 dB for the SQ scenario. Regarding the performance of speaker diarization, DER results increase when larger values for the weights $\alpha$ and $\beta$ of QQ and QS scenarios are used. These observations indicate that the model performance is sensitive to the loss weights under different scenarios. With the help of the multi-task interaction module, we can keep the performance for speaker extraction under SS and SQ scenarios and speaker diarization and simultaneously achieve a much lower power under the QQ scenario, which decreases from -1.93 dB/s to -24.00 dB/s.
V-G Complexity Analysis
In this section, we conduct a complexity analysis of the USED model in comparison to the baseline models, TS-VAD and SpEx+, as shown in Table VII. To ensure a fair comparison, the duration of the speech input for all models is kept consistent, with both the speech mixture and the speech references set to 4 seconds. The experimental results indicate that, compared to individual baseline models, the USED model has more parameters than the SpEx+ model but fewer than the TS-VAD model. However, the computational complexity of the USED model is higher than both SpEx+ and TS-VAD.
TABLE VIII: Impact of hyperparameter choices for embedding assignment module on model performance.
$m$ Metrics $p_{sel}$ $p_{em}$ $p_{res}$
0.1 0.3 0.5 0.7 0.9 0.1 0.3 0.5 0.7 0.9 0.1 0.3 0.5 0.7 0.9
1 SI-SDRi $\uparrow$ 11.81 10.36 12.12 12.81 12.60 10.77 12.12 12.93 13.09 13.10 13.12 12.99 12.72 13.01 12.12
DER $\downarrow$ 6.80 8.79 6.47 5.83 6.19 6.67 6.47 5.87 5.63 5.62 5.75 5.83 6.21 5.76 6.47
2 SI-SDRi $\uparrow$ 13.28 13.30 13.58 13.62 12.81 13.68 13.58 13.83 13.56 13.42 13.77 13.76 13.83 13.72 13.58
DER $\downarrow$ 5.17 5.11 4.98 5.28 9.41 5.01 4.98 5.32 4.97 5.05 5.67 5.15 5.24 4.93 4.98
all SI-SDRi $\uparrow$ 14.71 14.45 14.29 14.07 13.10 14.25 14.29 14.31 14.26 14.27 14.31 14.28 14.35 14.25 14.29
DER $\downarrow$ 3.87 4.00 4.22 4.41 6.27 4.30 4.22 4.16 4.13 4.18 4.23 4.22 4.16 4.19 4.22
TABLE IX: Model performance regarding different loss weight values on the validation set.
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}=0.1$ $\lambda_{3}=0.5$ $\lambda_{3}=1.0$
SI-SDRi $\uparrow$ DER $\downarrow$ SI-SDRi $\uparrow$ DER $\downarrow$ SI-SDRi $\uparrow$ DER $\downarrow$
0.1 0.1 14.31 4.15 14.27 4.26 14.27 4.15
0.1 0.5 14.01 3.85 13.99 3.76 14.06 3.74
0.1 1.0 13.54 3.85 13.46 4.00 13.32 4.19
0.5 0.1 14.22 5.01 14.13 5.14 14.19 5.01
0.5 0.5 14.16 4.23 14.27 4.26 14.32 4.14
0.5 1.0 14.19 3.97 14.32 3.92 14.28 3.84
1.0 0.1 14.11 5.49 14.15 5.37 14.17 5.47
1.0 0.5 14.15 4.57 14.18 4.55 14.32 4.45
1.0 1.0 14.22 4.21 14.32 4.14 14.29 4.22
It is important to note that the USED model simultaneously addresses two tasks: speaker extraction and diarization, whereas TS-VAD and SpEx+ are designed for a single task, with SpEx+ specifically extracting the speech for only one speaker. Therefore, we calculate the ‘Total MACs’. ‘Total MACs’ refers to the total number of MACs required for the models to process a speech mixture containing three speakers and obtain the corresponding speaker extraction and diarization results for all three speakers. When comparing the ‘Total MACs’, the ‘Total MACs’ of the USED and USED-F models are significantly lower than that of the baseline models combined. Specifically, the USED-F model reduces the total MACs by approximately 39%, while the USED model achieves a reduction of around 24%.
V-H Loss Weight Selection
To investigate the impact of different loss weights on the model’s performance and explore how to select appropriate values for these weights, we conduct a series of experiments with three loss weights ($\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$), as illustrated in Table IX. We evaluate the model’s performance on the validation set of the LibriMix dataset with max mode by testing each loss weight with values of 0.1, 0.5, and 1.0. The experimental results indicate that there may be room for improvement in the model’s performance regarding the selection of loss weights, although we set the values of all three weights to 1.0 for other experiments. For example, when $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are 0.5, 1.0 and 0.5, respectively, the USED model performs slightly better on both the speaker extraction and diarization tasks.
V-I Impact on Model Performance of Hyperparameter Tuning for Embedding Assignment Module
In this section, we examine the influence of hyperparameter selection within the embedding assignment module on the overall performance of the model, as shown in Table VIII. Specifically, we focus on three key hyperparameters associated with the embedding assignment module, conducting our analysis using the validation set from the LibriMix dataset in max mode. We assess the impact of varying the hyperparameter values $p_{sel}$, $p_{em}$, and $p_{res}$, which are assigned values of 0.1, 0.3, 0.5, 0.7, and 0.9, while keeping the two other hyperparameters fixed to their default value, across multiple configurations where $m$ is set to 1, 2, and all.
The experimental results demonstrate that the influence of the three hyperparameters on model performance varies depending on the value of $m$. For instance, for hyperparameter $p_{sel}$, larger values lead to better results when $m$ equals 1, whereas smaller values yield better performance when $m$ equals all. After carefully evaluating the overall performance across different values of $m$, we selected a set of parameters with $p_{sel}=0.5$, $p_{em}=0.3$, and $p_{res}=0.9$ for relatively good performance.
VI Conclusion
We propose USED, a unified network for universal speaker extraction and diarization, which integrates speaker extraction and diarization for managing speech mixtures with varying overlap ratios and variable number of speakers. We design an embedding assignment module to support producing results for a variable number of speakers based on speech references and enhance the model’s robustness. Additionally, a multi-task interaction module is designed to leverage information from both speaker extraction and diarization tasks. Experimental results demonstrate significant improvements in both highly and sparsely overlapped speech scenarios for the speaker extraction and speaker diarization tasks.